Please help me do this: 11-3(a+4)=3a-25
Answer by rothauserc(4718)

11 - 3(a+4) = 3a - 25
11 - 3a - 12 = 3a - 25
add 3a to both sides of =
11 - 12 = 6a - 25
-1 = 6a - 25
add 25 to both sides of =
6a = 24
a = 4
*********************************
check the answer by substituting a = 4 in the given equation
11 - 3(4+4) = 3*4 -25
-13 = -13
our answer checks